OpenStudy (anonymous):

Medal given factorize x^2 - 49

OpenStudy (anonymous):

x-7 x+7

OpenStudy (anonymous):

(x+7)(X-7)

OpenStudy (anonymous):

can you have 1(x^2-49)

OpenStudy (anonymous):

@cassytaylor92

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

1 is the factor of every expression, so we need not mention the '1' there. also, x^2-49 was still unfactorized there, right ?

OpenStudy (anonymous):

so, it would be right?

hartnn (hartnn):

for \(a^2-b^2 = (a+b)(a-b)\) so for \(x^2-7^2= (x+7)(x-7)\)

OpenStudy (anonymous):

so, it would not be right?

hartnn (hartnn):

so, 1(x^2-49) is not right and (x+7)(x-7) is correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so, how would you do, x^2-16=4x

hartnn (hartnn):

you want to solve that ? you know the quadratic formula ?

OpenStudy (anonymous):

quadratic formula

hartnn (hartnn):

you know what it is ?

OpenStudy (anonymous):

yes